OpenStudy (anonymous):

how can i found the x & y-intercept in this equation y - 6 = - 1/2 (x + 1). i need help pls

OpenStudy (anonymous):

Given eq is y - 6 = - 1/2 (x + 1) it can be written as 2 y - 12 = -x - 1 i.e. 2 y +x = 12 - 1 i.e. 2 y +x = 11 i.e \[\frac{2y}{11}+\frac{x}{11} =1 \rightarrow \frac{x}{11}+ \frac{2y}{11}=1 \] \[\frac{x}{11}+ \frac{y}{\frac{11}{2}}=1 \] comparing it with \[\frac{x}{a}+ \frac{y}{b}=1 \] we find x- intercept =a= 11 and y- intercept =b= 11/2